# Guided Multi-Fidelity Bayesian Optimization for Data-driven Controller Tuning with Digital Twins

### v1

@@ Abstract
We propose a guided multi-fidelity Bayesian optimization (GMFBO) framework for data-efficient controller tuning that integrates corrected digital twin (DT) simulations with real-world measurements. The method targets closed-loop systems with limited-fidelity simulations or inexpensive approximations. To address model mismatch, we build a multi-fidelity surrogate with a learned correction model that refines DT estimates from real data. An adaptive cost-aware acquisition function balances expected improvement, fidelity, and sampling cost. Our method ensures adaptability as new measurements arrive. The accuracy of DTs is re-estimated, dynamically adapting both cross-source correlations and the acquisition function. This ensures that accurate DTs are used more frequently, while inaccurate DTs are appropriately downweighted. Experiments on robotic drive hardware and supporting numerical studies demonstrate that our method enhances tuning efficiency compared to standard Bayesian optimization (BO) and multi-fidelity methods.
@@ I Introduction
In controller tuning, cost-effective simulation tools like Simulink, Gazebo, PyBullet, and LabVIEW have become essential, minimizing the need for expensive real-world experiments alongside digital twins [1]. Although digital twins often exhibit discrepancies when replicating real-world system behavior, they are nonetheless capable of capturing key operational dynamics and trends—making them valuable tools for informed optimization tasks [2]. Low-cost simulations, although less accurate, offer faster and cheaper evaluation cycles compared to high-fidelity experiments [3, 4]. For instance, they have been successfully employed in intelligent manufacturing to accelerate optimization processes [5, 6]. These simulations can enhance performance even at reduced fidelity [7], facilitating rapid prototyping and testing while decreasing reliance on costly physical experiments [8, 9].
@@
To leverage simulations with varying fidelities, Multi-Fidelity Bayesian Optimization (MFBO) has been developed, incorporating multiple information sources with different fidelity and cost levels [10, 11]. MFBO balances the trade-off between high-fidelity accuracy and low-fidelity efficiency, enabling more effective optimization strategies [12]. For example, [13] demonstrates MFBO in gait design for a soft quadruped, using simulation data alongside minimal physical trials to refine its controller. Advanced MFBO methods also account for input-dependent fidelities, where the discrepancy between low- and high-fidelity models varies with the decision variables. For instance, in robotics, a simulator may align well with the real system at nominal joint speeds (decision-variable settings) but deviate significantly at high accelerations. Such methods enhance data efficiency and robustness [14], underscoring the potential of integrating additional data to improve the precision and reliability of controller tuning [15].
@@
Recent works have explored online learning methods to improve control performance under changing system dynamics [16]. Optimization has advanced through correction models that align simulations with real-world data, enabling the creation of more accurate models [17]. These developments facilitate tuning frameworks where both controller and twin are updated during optimization, thereby enhancing data efficiency and robustness to changes in the real system [18, 19].
@@
The primary objective of this work is to improve the integration of digital twin data with real-system measurements, increase the data efficiency of BO-based controller tuning, and enable adaptivity to dynamic changes in the real system during optimization. We assume that data from the real system is expensive to acquire, while DT simulations are cheaper but of limited and variable accuracy. An important question we address is: How can DTs of varying accuracy help optimize system performance? To this end, our contributions are:
@@
  • A GMFBO framework for data-efficient controller tuning that integrates real and simulated data.
@@
  • A correction mechanism that refines DT predictions and quantifies their accuracy to guide optimization.
@@
  • A cost-aware acquisition function that balances fidelity, expected improvement, and sampling cost, mitigating reliance on inaccurate DTs.
  • Experimental validation on a robotic drive system, including experiments with changing real-system friction, demonstrating substantial improvements in tuning efficiency over baseline methods.
@@ II Controller Tuning Problem
Figure 1: Efficient controller parameters tuning based on multiple information sources
The target information source (IS1) is a closed-loop feedback system comprising the plant, controller, and associated components. We make no assumptions about the parametric controller and plant structures. Measurements from IS1 are accurate up to measurement noise. An additional information source (IS2) serves as the system’s digital twin. In many practical scenarios, the digital twin—constructed using physics-based modeling, empirical data, or a combination thereof—cannot be perfectly identified due to limited access to internal states, safety constraints, time limitations, or prohibitive data collection costs. Consequently, throughout the paper, we treat IS2 as less accurate than IS1. By accuracy (or fidelity), we mean the agreement between IS2’s predicted objective and IS1’s measured objective under identical controller parameters, formally defined in (12). Fig. 1 visualizes the problem setup.
Let $k\in\mathcal{K}$ be the controller parameter vector inside a known feasible set $\mathcal{K}\subseteq\mathbb{R}^{n_{k}}$. We choose the set $\mathcal{K}$ so that the step response of the target system and its digital twin are bounded in a fixed period $T$. We define $s\in[0,1]$ as a fidelity variable correlated with the accuracy of the information source. Here, $s=1$ denotes IS1 (highest fidelity), while smaller values indicate lower fidelity. Let $s^{\prime}<1$ denote the fidelity level associated with IS2. The augmented input random variable is $z:=[k,s]$ where $z\in\mathcal{K}\times\mathcal{S}$, with dimension $n_{z}:=n_{k}+1$. Let $\hat{\mathsf{g}}(z)\in\mathbb{R}$ such that $\hat{\mathsf{g}}(z):\mathcal{K}\times\mathcal{S}\to\mathbb{R}$ denote the information source–dependent objective function. There is no assumption on the convexity of the objective function.
We evaluate an information source $s$ at a given controller parameter $k$, measuring the feedback system output $\mathrm{y}(k,t,s)\in\mathbb{R}$ at time $t\in\mathbb{R}\subseteq[0,T]$. We aim to simultaneously optimize $n_{\mathrm{h}}>1$ performance metrics denoted by $h_{i}(\mathrm{y}(k,t,s))\in\mathbb{R}$ for $i=1,\ldots,n_{\mathrm{h}}$. These metrics quantify various aspects of system behavior, derived from measurable outputs relative to specified or desired references. We adopt a weighted sum approach to convert the multi-objective problem into a single-objective formulation [20], assigning relative importance to each performance metric [21]. So the overall performance cost function $\hat{f}(k)$ is
$\hat{\mathsf{g}}(k,s):=\mathrm{w}^{\mathsf{T}}\mathrm{h}(\mathrm{y}(k,t,s))=\sum_{i=1}^{n_{\mathrm{h}}}w_{i}h_{i}(\mathrm{y}(k,t,s)),\vskip-6.0pt$ (1)
where $\mathrm{h}(k):=[h_{1}(k),...,h_{n_{\mathrm{h}}}(k)]$ and $\mathrm{w}:=[w_{1},...,w_{n_{\mathrm{h}}}]$. The weights vector $\mathrm{w}$ adjusts the priority and scale of our metrics such that each metric properly contributes to the overall performance.
The data collection process is subject to measurement noise. Thus, a noisy observation $\mathsf{g}$ at input $z$ is modeled as
@@
where $\eta\sim\mathcal{N}(0,\sigma_{\eta}^{2})$ denotes zero-mean Gaussian noise with variance $\sigma_{\eta}^{2}$, estimated jointly with the kernel hyperparameters by maximizing the marginal log-likelihood [22]. This assumption is widely used because it effectively captures a broad range of independent random error sources, as per the central limit theorem [23]. We can imagine our information source including a performance assessment unit as a black-box, which provides the cost value $\mathsf{g}(z)$ given the controller parameter vector $k$. Now our data-driven controller tuning problem is to obtain the optimum controller parameter vector $k^{*}$ for IS1 by solving
$\begin{split}k^{*}:=\arg\min_{k\in\mathcal{K}}\ \hat{\mathsf{g}}(k,s=1).\end{split}\vskip-1.20007pt$ (3)
Evaluating the objective function $\hat{\mathsf{g}}(k,s=1)$ is expensive, since it involves operating the real setup.
@@ III Guided Multi-fidelity Bayesian Optimization
Figure 2: Overall schematic of our proposed guided multi-fidelity Bayesian optimization for tuning controller parameters
Fig. 2 depicts our GMFBO framework. We first describe the surrogate objective model used in BO. Next, after defining the new information source and estimating DT accuracy, we present the adaptive surrogate kernel and acquisition function, followed by the GMFBO algorithm.
@@ III-A Multi-fidelity Surrogate Model
We define a surrogate Gaussian process to model the multi-fidelity objective $\hat{\mathsf{g}}(z)$. We assume a GP prior over the function values $\mathbf{g}(\mathbf{z})\sim\mathrm{GP}(\mathbf{m}(\mathbf{z}),\mathbf{c}(z))$, where $\mathbf{g}(\mathbf{z}):=[\mathsf{g}(z_{1}),\dots,\mathsf{g}(z_{N_{0}+n})]^{\top}$ represents the vector of function values at specific training points $\mathbf{z}:=[z_{1},\dots,z_{N_{0}+n}]$, prior means $\mathbf{m}(\mathbf{z}):=[m(z_{1}),\dots,m(z_{N_{0}+n})]^{\mathsf{T}}$, prior covariances $\mathbf{c}(z):=[c(z,z_{1}),\dots,c(z,z_{N_{0}+n})]^{\mathsf{T}}$, $n$ is the number of observed points, $\mathbf{m}$ and $\mathbf{c}$ are the prior mean and covariance functions, respectively, $N_{0}:=N_{0_{s=1}}+N_{0_{s^{\prime}}}$ is the total size of initial data set
$\mathcal{D}_{\text{init}}:=\bigl\{\bigl(z_{i},\mathsf{g}(z_{i})\bigl)\>|\>i\in\{1,...,N_{0}\}\bigl\},\vskip-1.20007pt$ (4)
where $N_{0_{s=1}}$ is the initial IS1 data size and $N_{0_{s^{\prime}}}$ the initial IS2 data size. The likelihood function for $n$ observations is then $p(\mathbf{g}|\mathbf{z},\hat{\mathbf{g}})=\mathcal{N}(\hat{\mathbf{g}},\mathbf{\Sigma})$, where $\mathbf{\Sigma}=\text{diag}(\sigma_{\eta_{1}}^{2},\dots,\sigma_{\eta_{n}}^{2})$ is a diagonal matrix of observational noise variances. We define surrogate training dataset $\mathcal{D}\in\mathbb{R}^{(n+N_{0})\times(n_{z}+1)}$ with $n+N_{0}$ data points
$\mathcal{D}:=\bigl\{\bigl(z_{i},\mathsf{g}(z_{i})\bigl)\>|\>i\in\{1,...,n+N_{0}\}\bigl\}.$ (5)
The predictive distribution of $\hat{\mathsf{g}}(z_{*})$ at a test point $z_{*}$ is $\hat{\mathsf{g}}(z_{*})|\mathcal{D}\sim\mathcal{N}(\mu_{*},\sigma_{*}^{2})$, where the predictive mean and variance are respectively $\mu_{*}:=m(z_{*})+\mathbf{c}(z_{*})^{T}(\mathbf{C}+\mathbf{\Sigma})^{-1}(\mathbf{g}(\mathbf{z})-\mathbf{m}(\mathbf{z})),$ and $\sigma_{*}^{2}:=c(z_{*},z_{*})-\mathbf{c}(z_{*})^{T}(\mathbf{C}+\mathbf{\Sigma})^{-1}\mathbf{c}(z_{*}),$ with $\mathbf{c}(z_{*})$ being the covariance vector between $z_{*}$ and the training data, and $\mathbf{C}:=[c({z}_{\xi_{1}},{z}_{\xi_{2}})]\bigr\rvert_{\xi_{1},\xi_{2}=1}^{N_{0}+n}\in\mathbb{R}^{(N_{0}+n)\times(N_{0}+n)}$ [22].
@@ III-B Digital Twin Correction
Rather than modifying the objective function $\hat{\mathsf{g}}$, the proposed approach focuses on improving the digital twin’s state estimation through a learned correction. We propose a correction model, denoted by $\mathrm{GP}_{\text{c}}$, to refine the simulation output acquired from IS2 given the controller gains $k$, resulting in a corrected digital twin (IS3). This model is a Gaussian process regression model [22] trained to predict the output signal $\mathrm{y}(k,t,s=1)$ from IS1 using the estimated state $\mathrm{y}(k,t,s=s^{\prime})$ from IS2 at time $t$. The correction model is trained on correction dataset, a dataset of paired observations defined by
$\mathcal{D}_{\text{c}}:=\{([t,\mathrm{y}(k,t,s=s^{\prime})],\mathrm{y}(k,t,s=1))\mid t\in[0,T]\},\vskip-2.39996pt$ (6)
where $T$ is the period of data samples. The GP model assumes a homoscedastic Gaussian observation noise,
$\mathrm{y}(k,t,s=1)=\hat{\mathrm{y}}(k,t,s=1))+\epsilon,\vskip-2.39996pt$ (7)
where $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ is the output measurement noise with variances $\sigma_{\epsilon}^{2}$ which are jointly estimated with the kernel hyperparameters by maximizing the marginal log-likelihood [22]. Before training, both the inputs and targets are standardized to zero mean and unit variance, ensuring numerically stable hyperparameter learning [24].
@@
$\hat{\mathrm{y}}(k^{\prime},t,s=1):=\mu_{\text{c}}([t,\mathrm{y}(k^{\prime},t,s^{\prime})]).\vskip-2.39996pt$ (8)
The standard deviation of this prediction is then denoted by $\sigma_{\text{c}}([t,\mathrm{y}(k^{\prime},t,s^{\prime})])$. We define the prediction uncertainty as
$\bar{\sigma}_{\text{c}}:=\sqrt{\frac{1}{T}\sum_{t=0}^{T}\sigma^{2}_{\text{c}}([t,\mathrm{y}(k^{\prime},t,s^{\prime})])},\vskip-2.39996pt$ (9)
which measures the average predictive standard deviation of $\mathrm{GP}_{\text{c}}$ across the trajectory and reflects the reliability of the correction model. We impose the condition
$\bar{\sigma}_{\text{c}}\ll\alpha:=\sqrt{\frac{1}{T}\sum_{t=0}^{T}\mathrm{y}^{*}(t)}\in\mathbb{R},\vskip-2.39996pt$ (10)
with the threshold $\alpha\in\mathbb{R}$ serving as a normalized upper bound for acceptable uncertainty in IS3. If (10) holds, the model is deemed confident, and the data point is added as
$\mathcal{D}\leftarrow\mathcal{D}\cup\{[k^{\prime},s^{\prime}],\mathsf{g}_{\text{c}}(k^{\prime},s^{\prime})\}.\vskip-2.39996pt$ (11)
where $\mathsf{g}_{\text{c}}(k^{\prime},s^{\prime})$ is the performance estimated using IS3. For each IS1 query, this procedure is repeated to generate $n_{\text{c}}$ accepted IS3 samples that are added to $\mathcal{D}$. Example: For a unit step reference $\mathrm{y}^{*}(t)=1$ for $t\geq 0$, we obtain $\alpha=1$. Here, $\alpha$ equals $100\%$ of the normalized reference output, so $\bar{\sigma}_{\text{c}}\ll 1$ ensures the model’s uncertainty remains small relative to maximum expected magnitude.
Each time $\mathsf{g}_{\text{c}}(k^{\prime},s^{\prime})$ is estimated by IS3, the corresponding $\mathrm{y}(k^{\prime},t,s^{\prime})$ samples from IS2 are already available, so computing $\mathsf{g}_{\text{c}}(k^{\prime},s^{\prime})$ incurs no additional cost. We quantify the accuracy of IS2 relative to IS3 by
[…]
where $N_{\text{c}}$ is the total number of IS3 samples in $\mathcal{D}$.

### v2

@@ Abstract
We propose a guided multi-fidelity Bayesian optimization framework for data-efficient controller tuning that integrates corrected digital twin simulations with real-world measurements. The method targets closed-loop systems with limited-fidelity simulations or inexpensive approximations. To address model mismatch, we build a multi-fidelity surrogate with a learned correction model that refines digital twin estimates using real data. An adaptive cost-aware acquisition function balances expected improvement, fidelity, and sampling cost. Our method ensures adaptability as new measurements arrive. The digital twin accuracy is re-estimated, dynamically adapting both cross-source correlations and the acquisition function. This ensures that accurate simulations are used more frequently, while inaccurate simulation data are appropriately downweighted. Experiments on robotic drive hardware and supporting numerical studies demonstrate that our method enhances tuning efficiency compared to standard Bayesian optimization and multi-fidelity methods.
@@ I Introduction
In controller tuning, symbolic or physics-based simulation platforms reduce the reliance on costly real-world experiments[1]. Digital twins (DTs) can often capture key operational dynamics and trends, making them valuable tools for informed optimization tasks [2]. Low-cost simulations, although less accurate, offer faster and cheaper evaluation cycles compared to real experiments [3, 4]. For instance, they have been successfully employed in intelligent manufacturing to accelerate optimization processes [5, 6]. DTs can enhance performance even at reduced fidelity [7], facilitating rapid prototyping and testing while decreasing reliance on costly physical experiments [8, 9].
@@
Multi-Fidelity Bayesian Optimization (MFBO) has been developed to leverage simulations with varying levels of fidelity and cost[10, 11]. It incorporates multiple information sources[11]. MFBO balances the trade-off between high-fidelity accuracy and low-fidelity efficiency, enabling more effective optimization strategies [12]. For example, [13] demonstrates MFBO in gait design for a soft quadruped, using simulation data alongside minimal physical trials to refine its controller. Advanced MFBO methods also account for input-dependent fidelities, where the discrepancy between low- and high-fidelity models varies with the decision variables. For instance, in robotics, a simulator may align well with the real system at nominal joint speeds but deviate significantly at high accelerations. Such methods enhance data efficiency and robustness [14], underscoring the potential of integrating additional data to improve the precision and reliability of controller tuning [15].
@@
Recent works have explored online learning methods to improve control performance under changing system dynamics [16]. Optimization has advanced through correction models that align simulations with real-world data to create more accurate models [17]. These developments enable tuning frameworks where both the controller and a high-fidelity DT are updated during optimization, improving data efficiency and robustness to real-system changes [18, 19].
@@
We aim to improve the integration of low-fidelity, first-principles DTs with real-system measurements, increase the data efficiency of Bayesian optimization (BO) for controller tuning, and enable adaptivity to dynamic changes during optimization. We assume real-system data is costly, while DT simulations are cheaper but variably accurate, raising the question: How can imperfect DTs be exploited to optimize performance efficiently? Rather than seeking a globally accurate DT, we focus on selectively correcting it only in regions relevant to the optimization objective. To this end, our contributions are:
@@
  • A guided multi-fidelity Bayesian optimization (GMFBO) framework that integrates real and simulated data for data-efficient controller tuning.
@@
  • A correction mechanism that refines DT predictions where it matters for optimization and quantifies DT accuracy.
@@
  • A cost-aware acquisition function that balances fidelity, expected improvement, and sampling cost, mitigating overreliance on inaccurate simulations.
We validate GMFBO on a real robotic system, including scenarios with changing friction, demonstrating substantial gains in tuning efficiency over baseline methods.
@@ II Controller Tuning Problem
Figure 1: Efficient MFBO controller tuning using multiple information sources
We consider two information sources that provide data for controller tuning. The target information source (IS1) is a closed-loop system comprising the plant, controller, and associated components. We make no assumptions about the parametric controller and plant structures. Measurements from IS1 are accurate up to measurement noise. An additional information source (IS2) serves as the system’s digital twin, which is less accurate than IS1. By accuracy (or fidelity), we mean the agreement between IS2’s predicted objective and IS1’s measured objective under identical controller parameters. In many practical scenarios, the digital twin—constructed using physics-based modeling, empirical data, or a combination thereof—cannot be perfectly identified due to limited access to internal states or safety constraints.
Let $k\in\mathcal{K}$ be the controller parameter vector within a bounded box $\mathcal{K}\subseteq\mathbb{R}^{n_{k}}$. The set $\mathcal{K}$ is restricted to gains that ensure a stable, non-divergent unit step response on both the real system and the digital twin over the horizon $T$, thereby excluding unsafe controllers. We define $s\in\mathcal{S}=\{0,1\}$ as a binary fidelity variable correlated with source accuracy, where $s=1$ denotes the high-fidelity IS1 and $s=0$ the low-fidelity IS2. The augmented input random variable is $z:=[k,s]$ where $z\in\mathcal{K}\times\mathcal{S}$, with dimension $n_{k}+1$. Let $\hat{\mathsf{g}}:\mathcal{K}\times\mathcal{S}\to\mathbb{R}$ denote the information source–dependent objective function. There is no assumption on the convexity of the objective function.
We evaluate an information source $s$ at a given controller parameter $k$, measuring the feedback system output $\mathrm{y}(t,k,s)\in\mathbb{R}$, given the reference input signal $\mathrm{y}^{*}(t)\in\mathbb{R}$ at time $t\in[0,T]$. Performance metrics are $h_{i}(\mathrm{y}(t,k,s))\in\mathbb{R}$ for $i=1,\ldots,n_{\mathrm{h}}$. These metrics quantify various aspects of system behavior. We adopt a weighted sum approach to convert the multi-objective problem into a single-objective formulation [20], assigning relative importance to each performance metric [21]. So the overall performance cost function $\hat{f}(k)$ is
$\hat{\mathsf{g}}(k,s):=\mathrm{w}^{\mathsf{T}}\mathrm{h}(\mathrm{y}(t,k,s))=\sum_{i=1}^{n_{\mathrm{h}}}w_{i}h_{i}(\mathrm{y}(t,k,s)),\vskip-2.39996pt$ (1)
where $\mathrm{h}(\mathrm{y}):=[h_{1}(\mathrm{y}),...,h_{n_{\mathrm{h}}}(\mathrm{y})]$ and $\mathrm{w}:=[w_{1},...,w_{n_{\mathrm{h}}}]$. The weights vector $\mathrm{w}$ reflects the priority and scale of our metrics. Figure 1 visualizes the problem setup.
Due to measurement noise, an observation $\mathsf{g}$ at input $z$ is modeled as
@@
where $\eta\sim\mathcal{N}(0,\sigma_{\eta}^{2})$ denotes zero-mean Gaussian noise with variance $\sigma_{\eta}^{2}$. This assumption is widely used because it effectively captures a broad range of independent random error sources, as per the central limit theorem [22]. Our data-driven controller tuning problem is to obtain the optimum controller parameter vector $k^{*}$ for IS1 by solving
$\begin{split}k^{*}:=\arg\min_{k\in\mathcal{K}}\ \hat{\mathsf{g}}(k,1).\end{split}\vskip-2.39996pt$ (3)
Evaluating the objective function $\mathsf{g}(k,1)$ is expensive, since it involves operating the real setup.
@@ III Guided Multi-fidelity Bayesian Optimization
Figure 2 illustrates the information flow and interaction between the different sources in GMFBO. Data from IS1 is used both to correct the DT as an additional information source (IS3) and to provide high-fidelity observations for BO. The acquisition function selects the information source by incorporating the estimated DT mismatch (via the error metric $\hat{e}_{\text{IS2}}$) at each candidate. Most performance evaluations are delegated to the corrected DT in IS3, reducing the need for expensive IS1 queries. We next describe the surrogate objective model, the corrected source IS3 and its GP-based DT correction, the estimation of DT accuracy, the adaptive kernel and acquisition function, and finally the GMFBO algorithm.
Figure 2: Guided multi-fidelity Bayesian optimization for tuning controller parameters
@@ III-A Multi-fidelity Surrogate Model
We define a surrogate Gaussian process to model the multi-fidelity objective $\hat{\mathsf{g}}(z)$. A key contribution of our approach is to treat the fidelity variable $s$ as an explicit input to the GP, enabling a unified surrogate that seamlessly integrates data from all sources and implicitly weighs them according to their fidelity. We assume a GP prior over the function values $\mathbf{g}(\mathbf{z})\sim\mathrm{GP}(\mathbf{m}(\mathbf{z}),\mathbf{c}(z))$, where $\mathbf{g}(\mathbf{z}):=[\mathsf{g}(z_{1}),\dots,\mathsf{g}(z_{n})]^{\top}$ represents the vector of function values at specific training points $\mathbf{z}:=[z_{1},\dots,z_{n}]$, prior means $\mathbf{m}(\mathbf{z}):=[m(z_{1}),\dots,m(z_{n})]^{\mathsf{T}}$, prior covariances $\mathbf{c}(z):=[c(z,z_{1}),\dots,c(z,z_{n})]^{\mathsf{T}}$, $n$ is the number of observed points, $\mathbf{m}$ and $\mathbf{c}$ are the prior mean and covariance functions, respectively. Define initial data set
$\mathfrak{D}_{\text{init}}:=\bigl\{\bigl(z_{i},\mathsf{g}(z_{i})\bigl)\>|\>i\in\{1,...,N_{0_{s=1}}+N_{0_{s=0}}\}\bigl\},\vskip-2.39996pt$ (4)
where $N_{0_{s=1}}$ is the initial IS1 data size and $N_{0_{s=0}}$ the initial IS2 data size. The likelihood function for $n$ observations is then $p(\mathbf{g}|\mathbf{z},\hat{\mathbf{g}})=\mathcal{N}(\hat{\mathbf{g}},\mathbf{\Sigma})$, where $\mathbf{\Sigma}=\text{diag}(\sigma_{\eta_{1}}^{2},\dots,\sigma_{\eta_{n}}^{2})$ is a diagonal matrix of observational noise variances. The hyperparameters of the covariance kernel and observation noise are learned by maximizing the log marginal likelihood [23]. We define surrogate training dataset $\mathfrak{D}\in\mathbb{R}^{n\times(n_{k}+2)}$ with $n$ data points
$\mathfrak{D}:=\bigl\{\bigl(z_{i},\mathsf{g}(z_{i})\bigl)\>|\>i\in\{1,...,n\}\bigl\}.\vskip-2.39996pt$ (5)
The predictive distribution of $\hat{\mathsf{g}}(z_{*})$ at a test point $z_{*}$ is $\hat{\mathsf{g}}(z_{*})|\mathfrak{D}\sim\mathcal{N}(\mu_{*},\sigma_{*}^{2})$, where the predictive mean and variance are respectively $\mu_{*}:=m(z_{*})+\mathbf{c}(z_{*})^{T}(\mathbf{C}+\mathbf{\Sigma})^{-1}(\mathbf{g}(\mathbf{z})-\mathbf{m}(\mathbf{z})),$ and $\sigma_{*}^{2}:=c(z_{*},z_{*})-\mathbf{c}(z_{*})^{T}(\mathbf{C}+\mathbf{\Sigma})^{-1}\mathbf{c}(z_{*}),$ with $\mathbf{c}(z_{*})$ being the covariance vector between $z_{*}$ and the training data, and $\mathbf{C}:=[c({z}_{\xi_{1}},{z}_{\xi_{2}})]\bigr\rvert_{\xi_{1},\xi_{2}=1}^{n}\in\mathbb{R}^{n\times n}$ [23].
@@ III-B Digital Twin Correction
Our proposed approach focuses on improving the digital twin’s estimation of the performance cost function via a learned correction. The corrected digital twin $\mathrm{GP}_{\text{c}}$ is a Gaussian process regression model that predicts the real system output $\hat{y}(t,k,s=1)$ at each time step $t$ using the simulated output $y(t,k,s=0)$ from IS2 at that time step. $\mathrm{GP}_{\text{c}}$ is trained on a dataset of paired observations: $\mathfrak{D}_{\text{c}}:=\{([t,\mathrm{y}(t,k,s=0)],\mathrm{y}(t,k,s=1))\mid t\in[0,T]\}$, with Gaussian observation noise, $\mathrm{y}(t,k,s=1)=\hat{\mathrm{y}}(k,t,s=1)+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ is the output measurement noise. Before training, inputs and targets are normalized to zero mean and unit variance to support stable hyperparameter learning [24].
The correction model enables efficient data augmentation: each real evaluation at controller gains $k$ yields $n_{c}$ locally corrected samples around $k$. We set the total number of BO iterations on IS1 and IS2 as $N$. Each time the optimizer samples IS1 at controller gains $k$, we draw $k^{\prime}\sim\mathcal{N}(k,\,0.05\,\mathbb{K})\in\mathcal{K}$, where $\mathbb{K}$ is diagonal with entries equal to the feasible parameter ranges in $\mathcal{K}$. The factor $0.05$ sets the per-parameter standard deviation to $5\%$ of its range (elementwise, since $\mathbb{K}$ is diagonal), localizing IS3 evaluations near the latest IS1 point. We found that using a factor between 0.020.10.02-0.1 yields similar behavior. We draw $n_{\text{c}}$ corrected IS3 samples to the surrogate dataset $\mathfrak{D}$, per one real-system evaluation. If the draw $k^{\prime}$ falls outside $\mathcal{K}$, we project it back onto $\mathcal{K}$.
Then, IS2 generates a sequence of estimated outputs $\mathrm{y}(t,k^{\prime},0)$ for $t\in[0,\ldots,T]$. The sequence is processed point-by-point through $\mathrm{GP}_{\text{c}}$, and provides the corrected IS1 output at each time step, denoted by
$\mathrm{y}_{\text{c}}(t,k^{\prime},0):=\mu_{\text{c}}([t,\mathrm{y}(t,k^{\prime},0)]).\vskip-2.39996pt$ (6)
The standard deviation of this prediction is then denoted by $\sigma_{\text{c}}([t,\mathrm{y}(t,k^{\prime},0)])$. We define the prediction uncertainty as
$\bar{\sigma}_{\text{c}}:=\sqrt{\frac{1}{T}\sum_{t=0}^{T}\sigma^{2}_{\text{c}}([t,\mathrm{y}(t,k^{\prime},0)])},\vskip-2.39996pt$ (7)
which measures the average predictive standard deviation of $\mathrm{GP}_{\text{c}}$ across the trajectory and quantifies the confidence of the correction model in its predictions. We impose the condition
$\bar{\sigma}_{\text{c}}\ll\alpha:=\sqrt{\frac{1}{T}\sum_{t=0}^{T}(\mathrm{y}^{*}(t))^{2}}\in\mathbb{R},\vskip-2.39996pt$ (8)
where $\alpha$ is the root-mean-square (RMS) of the reference signal, providing a characteristic scale for evaluating prediction uncertainty. If (8) holds, the corrected prediction is deemed sufficiently confident, the data point is accepted and we update $\mathfrak{D}$: $\mathfrak{D}\leftarrow\mathfrak{D}\cup\{[k^{\prime},0],\mathsf{g}_{\text{c}}(k^{\prime},0)\}$, where $\mathsf{g}_{\text{c}}(k^{\prime},0)$ is the performance estimated using IS3. For each IS1 query, this procedure is repeated to generate $n_{\text{c}}$ accepted IS3 samples that are added to $\mathfrak{D}$. For a unit step reference $\mathrm{y}^{*}(t)=1$ for $t\geq 0$, we obtain $\alpha=1$. Here, $\alpha$ equals $100\%$ of the normalized reference output, so $\bar{\sigma}_{\text{c}}\ll 1$ ensures the model’s uncertainty remains small relative to maximum expected magnitude.
Adding several corrected points around each real measurement increases the local accuracy of the surrogate model, and reveals which directions in parameter space lead to improved performance. By providing this information, IS3 guides the optimizer towards promising regions of the objective landscape.
Each time $\mathsf{g}_{\text{c}}(k^{\prime},0)$ is estimated by IS3, the corresponding $\mathrm{y}(t,k^{\prime},0)$ samples from IS2 are already available, so computing $\mathsf{g}_{\text{c}}(k^{\prime},0)$ incurs no additional cost. We quantify the mismatch of IS2 relative to IS3 by
@@
where $N_{\text{c}}$ is the total number of IS3 samples in $\mathfrak{D}$. Large $\hat{e}_{\text{IS2}}$ indicates that IS2 predictions deviate from the corrected estimates, suggesting low simulation fidelity.